OpenStudy (anonymous):

solve 3b+4=16 for b A. b=12 B. b= 16 over 3 C. b= 4 D. b=13

OpenStudy (anonymous):

its b

OpenStudy (owlcoffee):

Whenever we see a single variable equation, in this case being the variable "b", it is read like this: "What number multiplied by 3 and added 4, gives me as a result, 16? Easy enough, all you would have to do is leave alone the variable on one side of the equality and that should be the value of it. So: \[3b+4=16\] if I sustract 4 to both sides, it'll look like this: \[3b=16-4\] and dividing by 4 on both sides, we get: \[b=\frac{ 16-4 }{ 3 }\] I'll leave the operating to you.

OpenStudy (anonymous):

3b+4=16 3b=16-4 3b=12 thus b=12/3 b=4

OpenStudy (anonymous):

answer is c